# Benchmark Dataset for Pore-Scale CO2-Water Interaction

### v1

@@ Abstract
Accurately capturing the complex interaction between CO2 and water in porous media at the pore scale is essential for various geoscience applications, including carbon capture and storage (CCS). We introduce a comprehensive dataset generated from high-fidelity numerical simulations to capture the intricate interaction between CO2 and water at the pore scale. The dataset consists of 624 2D samples, each of size $512\times 512$ with a resolution of $35\mu$m, covering 100 time steps under a constant CO2 injection rate. It includes various levels of heterogeneity, represented by different grain sizes with random variation in spacing, offering a robust testbed for developing predictive models. This dataset provides high-resolution temporal and spatial information crucial for benchmarking machine learning models.
@@
In this paper, we introduce a high-resolution dataset designed for benchmarking machine learning models in predicting CO2 behavior during multiphase flow in porous media. The dataset comprises 624 two-dimensional samples, each of size $512\times 512$ pixels with a spatial resolution of $35\mu$m, capturing the intricate interplay between CO2 and water over 100 equally spaced temporal snapshots under a constant CO2 injection rate. A distinctive feature of this dataset is its incorporation of varying levels of heterogeneity, represented through different grain sizes, which simulate realistic geological variability. This comprehensive dataset offers critical temporal and spatial granularity, serving as a utility for developing and benchmarking machine learning models.
@@ Methods
The geometries for the simulations are created from synthetic porous media of a single solid phase and initially the pore-space is filled with water. Initially, the original 1024 $\times$ 1024 domain was cropped into four subsamples. Each subsample was then flipped along the vertical axis, increasing the overall dataset size by a factor of eight. Each of these geometries is heterogeneous, i.e., the size of grains varies across the spatial domain (intra-sample heterogeneity). In addition, there are 5 levels of heterogeneity in the dataset (inter-sample heterogeneity) as shown in Figure 1. By exposing the ML models to a range of grain size distributions and spatial configurations, the dataset enhances the model’s ability to generalize to unseen porous media.
The inter-sample heterogeneity allows the ML model to develop robust feature extraction capabilities that are invariant to changes in grain sizes and configurations. This is crucial for ensuring that the predictions remain accurate across different geological formations. Similarly, the intra-sample heterogeneity forces the model to adapt to localized variations, which is essential for capturing fine-grained details in flow and saturation patterns. The dataset contains 624 geometries, each one is of size $512\times 512$ and the physical resolution per pixel is $35\mu$m. All samples are available in HDF5 format along with the simulations.
@@
The dataset has been made available on https://doi.org/10.5061/dryad.jm63xsjn5 and is organized into 10 folders, with each of the 5 geometries having its original version and a vertically flipped version ($2\times 5=10$). The simulation samples are provided in HDF5 format, with each file including water saturation ($\alpha_{water}$), pressure ($p$), capillary pressure ($pc$), horizontal velocity ($U_{x}$), vertical velocity ($U_{y}$), and a binary image of the physical domain (where pores are denoted by 1 and grains by 0), as detailed in Table 1 which also lists the keys required to access the data. The water saturation $\alpha_{water}$ is in the range $[0,1]$; hence, the CO2 saturation field can be computed using the relation $\alpha_{CO_{2}}=(1-\alpha_{water})\times\text{img}$, where img denotes the binary physical domain. Additionally, CSV files containing values for porosity, permeability, and relative permeability are provided, with details presented in Table 2.
@@
Usage Notes
The dataset is available in hdf5 format, which can be easily accessed and manipulated using Python libraries such as h5py. These files contain high-resolution data, including flow velocity components, pressure fields, and physical domain representations, making them suitable for scientific and engineering applications.

### v2

# A Benchmark Dataset for Machine Learning Surrogates of Pore-Scale CO2-Water Interaction

@@ Abstract
[…]
@@
[…]
@@ Methods
The pore structures are generated with the open‑source notebook DrawMicromodels.ipynb in https://github.com/hannahmenke/DrawMicromodels, commit 5e0f947, which perturbs a regular triangular lattice of mean grain radius $R_{0}$ by three heterogeneity amplitudes $\{\texttt{raddevmax},\texttt{xdevmax},\texttt{ydevmax}\}$. For the $n$-th grain
$R_{n}$ $=R_{0}\bigl(1+\delta^{(R)}_{n}\bigr)$,
$x_{n}$ $=x_{n}^{\text{lattice}}+L_{x}\,\delta^{(x)}_{n}$,
$y_{n}$ $=y_{n}^{\text{lattice}}+L_{y}\,\delta^{(y)}_{n}$,
where each perturbation term $\delta\in[-a,a]$ is sampled from a uniform distribution whose half‑width $a$ is the level‑dependent deviation listed in Table 1. Five levels are defined, ranging from well‑sorted media (Level 1) to highly heterogeneous media (Level 5).
Table 1: Quantitative definition of the five heterogeneity levels (dimensionless amplitudes).
Level 1 2 3 4 5
raddevmax 0.05 0.10 0.15 0.20 0.25
xdevmax 0.02 0.04 0.06 0.08 0.10
ydevmax 0.02 0.04 0.06 0.08 0.10
Physical motivation.
The radius variation mimics sedimentary sorting, while positional jitter reproduces local compaction and packing irregularities observed in outcrop sandstones ($C_{\!V}\approx 0.05$–0.25). Increasing these amplitudes therefore widens the pore‑throat distribution and the capillary contrast, both of which are known to control CO2–water displacement dynamics.
Parametric sweep and augmentation.
For each level we perform a deterministic sweep over $R_{0}\in\{70,80,90\}$ and target porosities $\phi\in\{0.20,0.25,0.30,0.35,0.40,0.45\}$, producing $5\times 3\times 6=90$ base images. Twelve images that displayed percolation shortcuts were discarded after visual inspection, leaving 78 accepted bases. Each $1024\times 1024$ image is subsequently cropped into four non‑overlapping quadrants ($512\times 512$), and mirrored vertically. This yields the final ensemble of $78\times 4\times 2=624$ geometries used in this study as shown in Figure 1. By exposing the ML models to a range of grain size distributions and spatial configurations, the dataset enhances the model’s ability to generalize to unseen porous media.
[…]
@@
The dataset has been made available on https://doi.org/10.5061/dryad.jm63xsjn5 [25] and is organized into 10 folders, with each of the 5 geometries having its original version and a vertically flipped version ($2\times 5=10$). The simulation samples are provided in HDF5 format, with each file including water saturation ($\alpha_{water}$), pressure ($p$), capillary pressure ($pc$), horizontal velocity ($U_{x}$), vertical velocity ($U_{y}$), and a binary image of the physical domain (where pores are denoted by 1 and grains by 0), as detailed in Table 2 which also lists the keys required to access the data. The water saturation $\alpha_{water}$ is in the range $[0,1]$; hence, the CO2 saturation field can be computed using the relation $\alpha_{CO_{2}}=(1-\alpha_{water})\times\text{img}$, where img denotes the binary physical domain. Additionally, CSV files containing values for porosity, permeability, and relative permeability are provided, with details presented in Table 3.
@@
To assess the dataset’s utility for improving model generalization, three models of a U-Net architecture [26] were trained on datasets of varying levels of heterogeneity. Each model was trained to predict future CO2 saturation by mapping a sequence of four consecutive saturation maps to the subsequent four timesteps. During evaluation, these models were applied in an autoregressive fashion to generate long-term predictions up to 60 timesteps. Model A was trained on the full dataset (5-Levels), model B was trained on a subset containing four of the five levels (4-Levels), and model C was trained on a subset with only the first level (1-Level). All models were then evaluated on samples from the fifth level, unseen by models B and C. For this analysis, all input samples were resized to $\times256$ pixels, and predictions were made for the first 60 timesteps.
The results, summarized in Table 4, indicate a clear benefit to training on a more diverse dataset. The 4-Levels model achieved a lower Mean Squared Error (MSE) on average (0.0254) across the test samples compared to the 1-Level model (0.0320). This demonstrates superior average performance and generalization. The 5-Levels model, having been trained on the test data, served as a benchmark and predictably achieved the lowest average MSE (0.0145). A direct visual comparison of the predicted simulations against the ground truth, as seen in Figure 7, corroborates these quantitative findings. Furthermore, the qualitative error maps in Figure 8 visualize this trend, showing progressively lower absolute error from the 1-Level to the 5-Levels model. However, the per-sample MSE plots in Figure 9 reveal that this improvement was not uniform across all samples; in some cases, the 4-Levels model performed similarly to, or slightly worse than, the 1-Level model. This suggests that while training on more varied data helps the model learn more general rules, it can also introduce biases that hinder performance on specific out-of-distribution samples. The primary conclusion is that increased training data diversity leads to better average generalization, though not necessarily universal improvement on every individual sample.
Table 4: Summary statistics for model performance on the unseen fifth level.
Model Name Mean MSE Final Step MSE Std Dev
5-Levels 0.014484 0.009853 0.004364
4-Levels 0.025410 0.023486 0.007635
1-Level 0.032036 0.037971 0.008166
Figure 7: Qualitative comparison of model predictions against the target simulation for a sample from the test set.
Figure 8: Prediction error maps for each model at different timesteps.
Figure 9: Mean Squared Error (MSE) over simulation timesteps for various samples of level 5.